OpenStudy (anonymous):

Factor and simplify the algebraic expression:

OpenStudy (anonymous):

\[x ^{5/6}-x ^{1/6}\]

OpenStudy (sanjanap):

\[6\sqrt{x ^{5}}\]

OpenStudy (sanjanap):

becasue, when you change that radical it becomes\[6\sqrt{x ^{5}} - 6\sqrt{x ^{1}}\] then, you simpify since they both have the smae radical. then, \[6\sqrt{x ^{5/1}}\] because, when it is - you divide. then the answer would be: \[6\sqrt{x ^{5}}\]

OpenStudy (sanjanap):

hope it helps :)

OpenStudy (anonymous):

Thank you! Do you do anything different with an equation like this : \[(x+9)^{1/4}+(x+9)^{3/4}\]

OpenStudy (sanjanap):

yes, use the same. Give em a medal, I ll solve it for you

OpenStudy (anonymous):

Okay